# Aligning Language Models to Explicitly Handle Ambiguity

### v2

@@ Abstract
In interactions between users and language model agents, user utterances frequently exhibit ellipsis (omission of words or phrases) or imprecision (lack of exactness) to prioritize efficiency. This can lead to varying interpretations of the same input based on different assumptions or background knowledge. It is thus crucial for agents to adeptly handle the inherent ambiguity in queries to ensure reliability. However, even state-of-the-art large language models (LLMs) still face challenges in such scenarios, primarily due to the following hurdles: (1) LLMs are not explicitly trained to deal with ambiguous utterances; (2) the degree of ambiguity perceived by the LLMs may vary depending on the possessed knowledge. To address these issues, we propose Alignment with Perceived Ambiguity (Apa), a novel pipeline that aligns LLMs to manage ambiguous queries by leveraging their own assessment of ambiguity (i.e., perceived ambiguity). Experimental results on question-answering datasets demonstrate that Apa empowers LLMs to explicitly detect and manage ambiguous queries while retaining the ability to answer clear questions. Furthermore, our finding proves that Apa excels beyond training with gold-standard labels, especially in out-of-distribution scenarios.
@@
Hyuhng Joon Kim1, Youna Kim1, Cheonbok Park2, Junyeob Kim1, Choonghyun Park1, Kang Min Yoo1 2 3, Sang-goo Lee1, Taeuk Kim4 * 1Seoul National University, 2NAVER Cloud 3NAVER AI LAB, 4Hanyang University {heyjoonkim, anna9812, juny116, pch330, sglee}@europa.snu.ac.kr {cbok.park, kangmin.yoo}@navercorp.com kimtaeuk@hanyang.ac.kr
@@ 2 Related Work
LLMs are typically trained through causal language modeling, a process essential for understanding and generating text of high fluency and consistency. To better harness these models, approaches have been developed to align them with human preferences (Leike et al., 2018; Ji et al., 2023b) through various forms, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Chakraborty et al., 2024), and Supervised Fine-tuning (SFT) (Dong et al., 2023; Yang et al., 2023; Zhou et al., 2024). Previous works focused on preferences such as helpfulness (Ding et al., 2023; Köpf et al., 2023; Xu et al., 2024), safety (Bai et al., 2022; Ji et al., 2023a; Liu et al., 2024b), and factuality (Yang et al., 2023; Tian et al., 2024). Building on this foundation, our research expands the scope of research by focusing on aligning models to understand and manage ambiguity effectively.
@@ 3.2 Perceived Ambiguity Detection
This stage aims to identify samples from that the model perceives as ambiguous. Given that it is challenging for the model to express ambiguity explicitly, we construct a proxy task to estimate the ambiguity from the model’s perspective. Specifically, the model is prompted to self-disambiguate the given query and generate a disambiguation . The model leverages its intrinsic knowledge related to to generate further details in this process. If is underspecified and the model possesses related knowledge necessary to compensate, then would yield a higher certainty (lower entropy) from the model’s perspective. On the other hand, if requires no specification or the model lacks the necessary knowledge, would exhibit a similar level of uncertainty as . To quantify the uncertainty associated with and , we employ the model’s average entropy (Malinin and Gales, 2021; Abdar et al., 2021). Formally, the entropy of an output distribution is defined as follows:
@@ 3.3 Response Construction
We utilize a pre-defined clarification request as for .

### v3

@@ Abstract
In interactions between users and language model agents, user utterances frequently exhibit ellipsis (omission of words or phrases) or imprecision (lack of exactness) to prioritize efficiency. This can lead to varying interpretations of the same input based on different assumptions or background knowledge. It is thus crucial for agents to adeptly handle the inherent ambiguity in queries to ensure reliability. However, even state-of-the-art large language models (LLMs) still face challenges in such scenarios, primarily due to the following hurdles: (1) LLMs are not explicitly trained to deal with ambiguous utterances; (2) the degree of ambiguity perceived by the LLMs may vary depending on the possessed knowledge. To address these issues, we propose Alignment with Perceived Ambiguity (Apa), a novel pipeline that aligns LLMs to manage ambiguous queries by leveraging their own assessment of ambiguity (i.e., perceived ambiguity). Experimental results on question-answering datasets demonstrate that Apa empowers LLMs to explicitly detect and manage ambiguous queries while retaining the ability to answer clear questions. Furthermore, our finding proves that Apa excels beyond training with gold-standard labels, especially in out-of-distribution scenarios. The data and code are available at https://github.com/heyjoonkim/APA.
@@
Hyuhng Joon Kim1, Youna Kim1, Cheonbok Park2 3, Junyeob Kim1, Choonghyun Park1, Kang Min Yoo1 2 4, Sang-goo Lee1 5, Taeuk Kim6 * 1Seoul National University, 2NAVER Cloud, 3KAIST AI, 4NAVER AI LAB, 5IntelliSys, Korea, 6Hanyang University {heyjoonkim, anna9812, juny116, pch330, sglee}@europa.snu.ac.kr {cbok.park, kangmin.yoo}@navercorp.com kimtaeuk@hanyang.ac.kr
@@ 1 Introduction
Our contributions can be summarized as follows:
1.
We propose a novel approach, Alignment with Perceived Ambiguity (Apa), which enables language models to explicitly handle ambiguous inputs by leveraging perceived ambiguity.
2.
We introduce three new datasets— AmbigTriviaQA, AmbigWebQuestions, and AmbigFreebaseQA—specifically designed to evaluate the model’s capability of addressing ambiguity.
3.
Through empirical validation on multiple question-answering datasets, we demonstrate that Apa enables models to effectively handle ambiguous queries.
@@ 2 Related Work
LLMs are typically trained through causal language modeling, a process essential for understanding and generating text of high fluency and consistency. To better harness these models, approaches have been developed to align them with human preferences (Leike et al., 2018; Ji et al., 2023b). This has taken various forms, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Chakraborty et al., 2024), and Supervised Fine-tuning (SFT) (Dong et al., 2023; Yang et al., 2023; Zhou et al., 2024). Previous works focused on preferences such as helpfulness (Ding et al., 2023; Köpf et al., 2023; Xu et al., 2024), safety (Bai et al., 2022; Ji et al., 2023a; Liu et al., 2024b), and factuality (Yang et al., 2023; Tian et al., 2024). Building on this foundation, our research expands the scope by focusing on aligning models to understand and manage ambiguity effectively.
@@ 3.2 Perceived Ambiguity Detection
This stage aims to identify samples from that the model perceives as ambiguous. Given that it is challenging for the model to express ambiguity explicitly, we construct a proxy task to estimate the ambiguity from the model’s perspective. Specifically, the model is prompted to self-disambiguate the given query and generate a disambiguation . The model leverages its intrinsic knowledge related to to generate further details in this process. If is underspecified and the model possesses related knowledge necessary to compensate, then would yield a higher certainty from the model’s perspective. On the other hand, if requires no specification or the model lacks the necessary knowledge, would exhibit a similar level of uncertainty as . To quantify the uncertainty associated with and , we employ the model’s average entropy (Malinin and Gales, 2021; Abdar et al., 2021). Formally, the entropy of an output distribution is defined as follows:
@@ 3.3 Response Construction
We utilize a pre-defined clarification request as for . Specifically, a list of clarification requests is pre-defined, and a single response is randomly selected as for each instance.
@@ 7 Conclusion
Acknowledgement
This work was partly supported by SNU-NAVER Hyperscale AI Center and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2020-II201373, Artificial Intelligence Graduate School Program (Hanyang University), NO.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]